Shubham painted 2 / 3 of the wall space in his room. His sister Madhavi helped and painted 1 / 3 of the wall space. How much did they paint together?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Wall space painted by Shubham in a room = 2 / 3
Wall space painted by Madhavi in a room = 1 / 3
Total space painted by both = (2 / 3 + 1 / 3)
= (2 + 1) / 3
= 3 / 3
= 1
∴Shubham and Madhavi together painted 1 complete wall in a room.

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Fraction of wall painted by Shubham = 2/3
Fraction of wall painted by Madhavi = 1/3
Total fraction of wall painted = 2/3+1/3 = (2+1)/3 = 3/3= 1
 
Therefore, Together they painted 1 wall in a room.
 
Hope it helps!
Thank you